# Second- and third-order properties of multidimensional Langevin equations

### v13

@@ 6.8. Quantitative significance of third-order effects
The above applies only to stationary variables. For integrated variables, we compare the quadratic variations of the two terms in Eq. (177). We can easily generalize this to two integrated variables $y$ and $z$. We need the quantity:
(199) ${\widetilde{M}}^{iji^{\prime}j^{\prime}}:=\left\langle{\frac{\mathrm{d}[x^{i}x^{j},x^{i^{\prime}}x^{j^{\prime}}]}{\mathrm{d}t}}\right\rangle=2\left({C^{ii^{\prime}}D^{jj^{\prime}}+C^{ij^{\prime}}D^{ji^{\prime}}+C^{ji^{\prime}}D^{ij^{\prime}}+C^{jj^{\prime}}D^{ii^{\prime}}}\right).$
Similarly to before, we need the “symmetric inverse” $\widetilde{N}$ of the above, neglecting permutations:
(200) ${\widetilde{M}}^{iji^{\prime}j^{\prime}}{\widetilde{N}}_{i^{\prime}j^{\prime}kl}=\frac{\delta^{i}_{k}\delta^{j}_{l}+\delta^{i}_{l}\delta^{j}_{k}}{2}.$
This results in the following “ensemble covariance”:
@@
where the expectation is evaluated in the linear model. It is seen that the above satisfies Eq. (93). We then have the “ensemble covariance” of these quantities collectively:
@@
As for second-order quantities, if the time-step is not small, it may be suitable to use discrete-time estimators of the diffusion.
@@ 6.10. Comparison of experimental data and theoretical predictions
We consider the sum and difference of these quantities at $\tau=0$. The former is equal to $(\Delta t)L(x^{i}x^{j},y)$, while the latter is equal to $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. We can calculate the “ensemble covariances” of these quantities using the aforementioned prescriptions. We may consider using them for comparison at arbitrary $\tau$. However, we may be interested in comparing the above quantities directly. They may be obtained by taking one-half the sum and difference of the sum and difference. We may treat the sum and difference of the above quantities as uncorrelated in the “ensemble covariance” as they acquire opposite signs under time reversal. Thus we conclude that the above quantities should be compared according to one-fourth the sum of the “ensemble covariances” of $(\Delta t)L(x^{i}x^{j},y)$ and $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. As in the linear case, and as is analogously true for all quantities, if the theoretical values at $\tau=0$ exceed the prescriptions of the “ensemble variances” (for either second- or third-order quantities), we may need to increase the tolerance for comparison accordingly.
@@ 7.1. Linear Gaussian systems
So far, we have considered variables that are even under time reversal. However, some so-called “odd” variables, like velocity, change sign upon time reversal. We first consider linear Gaussian systems. Our state variable contains some even variables, denoted $\mathbf{x}^{+}$, and some odd variables, denoted $\mathbf{x}^{-}$. The dynamical matrix $\mathbf{A}$ can be partitioned into various components depending on which variables they connect, viz.:
@@ 8.2. Almost Markovian dynamics
We see that the condition $\mu^{2}\gg 1$ amounts to separation of time-scales. More specifically, the squares of the real parts of the eigenvalues of $\mu\mathbf{A}_{\mathbf{v}}$ should be large compared to the complex moduli of the eigenvalues of $\mathbf{A}_{\mathbf{x}}$. We also assume that the real parts of the eigenvalues of $\mathbf{A}_{\mathbf{x}}$ are at least of order 1 in magnitude. The inverse transformation is given by:
@@ 8.3. Third-order properties
We now consider dynamics involving inhomogeneous diffusion and nonlinearity, as follows:
@@
where all coefficients are constant. We first note that the estimation of quantities $\langle f(\mathbf{x},\mathbf{v})\dot{\mathbf{v}}\rangle$ from measured time-series data $\mathbf{x}(n\Delta t)$, $n=0,1,2,\ldots$ is not what might be naïvely expected from time-discretization of $\mathbf{v}$ and $\dot{\mathbf{v}}$, but has a correction due to fluctuations [10]. Aside from that, there is the issue of measuring quantities that are robust against measurement noise, for which the authors of Ref. [10] consider quantities of the form:

### v14

@@ 6.8. Quantitative significance of third-order effects
The r.h.s. is the number of degrees of freedom of $a^{i}_{jk}$. The above applies only to stationary variables. Alternatively, we may compare the quadratic variations of the linear and quadratic terms in the drift function. We thus introduce the quantity:
(199) ${\widetilde{M}}^{jkj^{\prime}k^{\prime}}:=\left\langle{\frac{\mathrm{d}[x^{j}x^{k},x^{j^{\prime}}x^{k^{\prime}}]}{\mathrm{d}t}}\right\rangle=2\left({C^{jj^{\prime}}D^{kk^{\prime}}+C^{jk^{\prime}}D^{kj^{\prime}}+C^{kj^{\prime}}D^{jk^{\prime}}+C^{kk^{\prime}}D^{jj^{\prime}}}\right),$
evaluated in the linear Gaussian model. Similarly to before, we need the “symmetric inverse” $\widetilde{N}$ of the above, neglecting permutations:
(200) ${\widetilde{M}}^{jkj^{\prime}k^{\prime}}{\widetilde{N}}_{j^{\prime}k^{\prime}lm}=\frac{\delta^{j}_{l}\delta^{k}_{m}+\delta^{j}_{m}\delta^{k}_{l}}{2}.$
We then specify:
(201) $a^{i}_{jk}a^{i^{\prime}}_{j^{\prime}k^{\prime}}\sim 2A^{i}_{l}A^{i^{\prime}}_{l^{\prime}}D^{ll^{\prime}}{\widetilde{N}}_{jkj^{\prime}k^{\prime}}.$
The advantage of this formulation is that it can be extended to integrated variables. For a description independent of linear transformations of coordinates, we compare the linear and quadratic terms in Eq. (177). For two integrated variables $y$ and $z$, we have the “ensemble covariances”:
(202) $a^{i}_{jk}\left({a^{y}_{i^{\prime}j^{\prime}}-A^{y}_{k^{\prime}}(\mathbf{A}^{-1})^{k^{\prime}}_{l^{\prime}}a^{l^{\prime}}_{i^{\prime}j^{\prime}}}\right)\sim 2{\widetilde{N}}_{jkk^{\prime}l^{\prime}}A^{i}_{l}(A^{k^{\prime}}_{i^{\prime}}\delta^{l^{\prime}}_{j^{\prime}}+\delta^{k^{\prime}}_{i^{\prime}}A^{l^{\prime}}_{j^{\prime}})\left({D^{ly}-A^{y}_{p}(\mathbf{A}^{-1})^{p}_{q}D^{lq}}\right),$
@@
where the expectation is evaluated in the linear Gaussian model. It is seen that the above satisfies Eq. (93). We then have the “ensemble covariance” of these quantities collectively:
@@ 6.10. Comparison of experimental data and theoretical predictions
We consider the sum and difference of these quantities at $\tau=0$. The former is equal to $(\Delta t)L(x^{i}x^{j},y)$, while the latter is equal to $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. We can calculate the “ensemble variances” of these quantities using the aforementioned prescriptions. We may consider using them for comparison at arbitrary $\tau$. However, we may be interested in comparing the above quantities directly. They may be obtained by taking one-half the sum and difference of the sum and difference. We may treat the sum and difference of the above quantities as uncorrelated in the “ensemble covariance” as they acquire opposite signs under time reversal. Thus we conclude that the above quantities should be compared according to one-fourth the sum of the “ensemble variances” of $(\Delta t)L(x^{i}x^{j},y)$ and $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. As in the linear case, and as is analogously true for all quantities, if the theoretical values at $\tau=0$ exceed the prescriptions of the “ensemble variances” (for either second- or third-order quantities), we may need to increase the tolerance for comparison accordingly.
@@ 7. Odd variables
So far, we have considered variables that are even under time reversal. However, some so-called “odd” variables, like velocity, change sign upon time reversal. In this section, we will discuss the conditions for time-reversal symmetry in the presence of odd variables, to third order.
@@ 7.1. Linear Gaussian systems
We first consider linear Gaussian systems. Our state variable contains some even variables, denoted $\mathbf{x}^{+}$, and some odd variables, denoted $\mathbf{x}^{-}$. The dynamical matrix $\mathbf{A}$ can be partitioned into various components depending on which variables they connect, viz.:
@@ 8.2. Almost Markovian dynamics
We see that the condition $\mu^{2}\gg 1$ amounts to separation of time-scales. More specifically, the squares of the real parts of the eigenvalues of $\mu\mathbf{A}_{\mathbf{v}}$ should be large compared to the complex moduli of the eigenvalues of $\mu^{-1}\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{A}_{\mathbf{x}}$. We shall assume that the real parts of the eigenvalues of $\mathbf{A}_{\mathbf{v}}$ and of $\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{A}_{\mathbf{x}}$ are at least of order 1 in magnitude. The inverse transformation is given by:
@@
We may evaluate quantitative significance of the velocity-velocity angular momenta as:
(325) $L(v^{i},v^{j})L(v^{k},v^{l})\sim 2(D^{ik}D^{jl}-D^{il}D^{jk}).$
@@
Finally, we address position-velocity covariance functions with a time-lag. For $\tau\gg\mathcal{O}(\mu^{-1})$, we have:
(330) $\langle v^{i}(\tau)x^{j}(0)\rangle=\langle\dot{\widetilde{x}}^{i}(\tau)\widetilde{x}^{j}(0)\rangle+\mathcal{O}(\mu^{-3}),$
which suggests the criterion:
(331) $\Delta_{\textrm{exp\textendash theo}}\langle v^{i}(\tau)x^{j}(0)\rangle\Delta_{\textrm{exp\textendash theo}}\langle v^{k}(\tau)x^{l}(0)\rangle\sim\mu^{-2}\widetilde{D}^{ik}\widetilde{D}^{jl}$
where $\widetilde{\mathbf{D}}:=\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}$ and, as before, $\Delta_{\textrm{exp\textendash theo}}$ denotes the deviation of experimental to theoretical values. However, extrapolating the above to $\tau=0$ and antisymmetrizing results in one-fourth the prescription for position-velocity covariances at zero time-lag. This suggests an interpolation of the form:
(332) $\Delta_{\textrm{exp\textendash theo}}\langle v^{i}(\tau)x^{j}(0)\rangle\Delta_{\textrm{exp\textendash theo}}\langle v^{k}(\tau)x^{l}(0)\rangle\sim\mu^{-2}(\delta^{i}_{i^{\prime}}+\exp(\mu\mathbf{A}_{\mathbf{v}}\tau)^{i}_{i^{\prime}})(\delta^{k}_{k^{\prime}}+\exp(\mu\mathbf{A}_{\mathbf{v}}\tau)^{k}_{k^{\prime}})\widetilde{D}^{i^{\prime}k^{\prime}}\widetilde{D}^{jl}.$
As before, if $\langle\mathbf{x}\mathbf{v}^{\mathsf{T}}\rangle$ or $\mathbf{L}(\mathbf{v},\mathbf{v}^{\mathsf{T}})$ exceed the “reference” values, this needs to be accounted for in the above comparison. However, this prescription may be somewhat problematic if $\mathbf{A}_{\mathbf{v}}$ has complex eigenvalues. In such a case, the r.h.s. of the above may be oscillatory and could be close to zero for $\tau\sim\mathcal{O}(\mu^{-1})$ [^20] We might consider ways to eliminate the complex eigenvalues. One idea is to use the time-symmetric quantity $\mathbf{D}\langle\mathbf{v}\mathbf{v}^{\mathsf{T}}\rangle^{-1}$ in place of $\mathbf{A}_{\mathbf{v}}$. However, this will in general change the real parts of the eigenvalues (see remarks at the end of section 2.1). Another idea is to use the eigendecomposition of $\mathbf{A}_{\mathbf{v}}$ and simply change all the eigenvalues to their real parts. However, this mapping cannot be continuously extended to non-diagonalizable matrices. To see this, consider a matrix with eigenvalues $\lambda$ and $\lambda+\delta\lambda$ with corresponding eigenvectors $\mathbf{w}$ and $\mathbf{w}+\delta\mathbf{w}$, where $\delta\lambda\neq 0$, $\delta\mathbf{w}$ are infinitesimal. Then $\delta\mathbf{w}/\delta\lambda$ is a generalized eigenvector corresponding to $\mathbf{w}$. Thus the matrix in question depends only on $\delta\mathbf{w}/\delta\lambda$, and not on the complex argument of $\delta\lambda$. However, the proposed procedure will have different results depending on whether $\delta\lambda$ is real or imaginary, and thus it cannot be continuously extended to this case..
@@ 8.3. Third-order properties
We now consider dynamics involving inhomogeneous diffusion and nonlinear drift, as follows:
@@
where all coefficients are constant.
8.3.1. Inference from data
We first note that the estimation of quantities $\langle f(\mathbf{x},\mathbf{v})\dot{\mathbf{v}}\rangle$ from measured time-series data $\mathbf{x}(n\Delta t)$, $n=0,1,2,\ldots$ is not what might be naïvely expected from time-discretization of $\mathbf{v}$ and $\dot{\mathbf{v}}$, but has a correction due to fluctuations [10]. Aside from that, there is the issue of measuring quantities that are robust against measurement noise, for which the authors of Ref. [10] consider quantities of the form: